Example usage of the todo package

Sven G. Robertz

1 Foo

Blah, blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah

ToDo:

More blah blah here?

blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah

2 Bar

Blah, blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah blah

Yada yada yada yada yada yada yada yada yada yada yada yada yada yada yada yada yada yada yada yada yada

ToDo

□ Section 1, page 1: More blah blah here? √ Section 2, page 1: Write some yada yada Is this enough?